OpenStudy (anonymous):

Graph the solution for the following linear inequality system. Click on the graph until the final result is displayed. y > 5 y ≥ x

OpenStudy (anonymous):

Solve the following system of equations by the substitution method. 5x = y + 6 2x - 3y = 4 What is the value of the y-coordinate?

OpenStudy (anonymous):

y = Y Y-INTERCEPT (0,5)

OpenStudy (anonymous):

Given the system of equations, what is the value of the system determinant? 2x + y = 8 x - y = 10

OpenStudy (anonymous):

6,4

OpenStudy (anonymous):

Given the system of equations, what is the value of the system determinant? 2x + y = 8 x - y = 10 -1 -2 -3